OpenStudy (kitkat1):

Part A: If (6^2)^x = 1, what is the value of x? Explain your answer. Part B: If (6^0)^x = 1, what are the possible values of x? Explain your answer.

OpenStudy (aum):

(6^2)^x = 1 36^x = 1 x = 0 because 36^0 = 1

OpenStudy (kitkat1):

What is the answer for the second one?

OpenStudy (aum):

What is 6^0 equal to?

OpenStudy (kitkat1):

1?

OpenStudy (kitkat1):

so its a 0 to?

OpenStudy (aum):

(6^0)^x = 1 (1)^x = 1 The right hand side is 1 which can be written either as 1^0 or as 1^1 both of which are 1. 1^x = 1^0 implies x = 0 or 1^x = 1^1 implies x = 1 x = 0 or 1.

OpenStudy (kitkat1):

So 0 or 1 would work?

geerky42 (geerky42):

Hmm... \(1^2=1\cdot1=1\\~\\1^3=1\cdot1\cdot1=1\\~\\1^4=1\cdot1\cdot1\cdot1=1\\~\\1^5=1\cdot1\cdot1\cdot1\cdot1=1\) Apparently it works for any number.

OpenStudy (kitkat1):

Thanks